OpenStudy (anonymous):

The volume of a pyramid with a square base is 32 cubic centimeters. If its height is 4 centimeters more than the side of its base, what is the height? 2 centimeters 4 centimeters 6 centimeters 8 centimeters

OpenStudy (p0sitr0n):

\[V=1/3* h * A_{base}\] but h=a+4, A=(h-4)^2 now you plug that back in, \[V=1/3* h * (h-4)^2\] now you can solve for h,since you know V

OpenStudy (p0sitr0n):

if you do not want to solve this equation, just plug in all of the choice you have for h, and check whether the equality holds

OpenStudy (anonymous):

Is it 6? :) @P0sitr0n